OpenStudy (anonymous):

find the perpendicular distance between y=2x+2 and y=2x+7

OpenStudy (abb0t):

Well, as you can see, the slope of both of these is \(\sf \color{red}{2}\), which means the perpendicular is \(\sf \color{red}{-\frac{1}{2}}\). That means, get "x" and "y" points. the intersection with the first line is where \(\sf \color{blue}{-\frac{1}{2} x + 2 = 2x + 7}\) x = -2 then y = 2(-2)+7 = 3 So your first point is (-2,3). Now, can you find your second? You can pick any point on the line. Then use distance formula to get the distance.